Evaluate log9 9 exponent of 3
Answer by stanbon(75887)

Evaluate log9(9)^3
----------------------
Let log9(9^3) = x
---
base = 9
log = exponent = x
result = 9^3
----
Write in exponential form:
9^x = 9^3
---
x = 3
=======
So, log9(9^3) = 3
=====================
Cheers,
Stan H.
=============